# TaoSR1: The Thinking Model for E-commerce Relevance Search

### v3

@@ 4.1.1. Dataset and Metrics
4.1.2. Implementation Details
During the Supervised Fine-Tuning (SFT) phase, we used a batch size of $1024$ and a learning rate of 1e61e-6, managed by a cosine scheduler with a warmup ratio of $0.05$. We employed the AdamW (loshchilov2018decoupled, ) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model was trained for $2$ epochs with a maximum input length of $4096$ tokens.
In the GRPO reinforcement learning phase, we configured a rollout batch size of $64$ with a learning rate of 1e61e-6. For each sample, we generated $16$ responses using a sampling temperature of $0.99$ and top-k of $100$. The thresholds for our online difficulty sampling were set to a range of $[0.01,0.9]$. For stabilization, we clipped the importance sampling ratio to a range of $1\pm 0.2$ and the advantage values to $\pm 2$. Our training pipeline is implemented based on the open-source framework ROLL(wang2025reinforcementlearningoptimizationlargescale, ).
@@ 4.1.3. Baselines
We employ multiple models as our baselines, including BERT, Qwen3-0.6B, and Qwen3-30BA3B(yang2025qwen3technicalreport, ). Our model is iteratively developed based on Tbstar. Tbstar is a closed-source e-commerce large language model foundation trained from scratch by Taobao, utilizing over 1010T tokens of general corpus and rich domain-specific e-commerce corpus. Based on a Mixture of Experts (MoE) architecture, it has a total parameter count of 4242B with 3.53.5B activated parameters.
  • BERT: A 24-layer BERT model foundation trained from scratch based on general corpus and Taobao’s internal e-commerce corpus, subsequently fine-tuned on the full training set using the formula 4.
@@
  • Qwen3-30BA3B: An open-source Mixture-of-Experts (MoE) model from the Qwen family, featuring 30 billion total parameters and 3 billion active parameters per forward pass. This model was also supervised fine-tuned (SFT) on our full training set same as Qwen3-0.6B.
@@ 4.3.1. Ablation of Post-CoT
We contrast the ’think-then-respond’ paradigm with the ’respond-then-think’ paradigm (distinguished by the postCoT subscript) in Table 2. We evaluate the performance of each paradigm both at the SFT stage and subsequent DPO stage. At the SFT-only stage, the post-CoT approach demonstrates a significant advantage over the standard approach. This provides strong evidence that the post-CoT effectively mitigates the issue of error accumulation in CoT, positioning it as a potentially superior approach for classification tasks. Furthermore, the application of DPO consistently yields notable performance gains under both configurations. This confirms that even within the post-CoT framework, there remains substantial headroom for improvement through reinforcement learning techniques.
@@ 4.3.3. Ablation of CumPT
This section provides a detailed analysis of the advantages of our proposed CumPT method. Macro F1 is computed directly on the model’s raw outputs, representing its intrinsic classification capability. Online Macro F1 is computed after the tiering mentioned in Sec. 3.4 (either the traditional formula-based approach via Eq.11, Eq.12 or CumPT). Additionally, we report the Precision, Recall, and F1-score specifically for the “Good” tier.
To illustrate the sensitivity of the traditional method, we fixed a subset of its many hyperparameters ($\alpha_{1}=0.33$, $\alpha_{2}=0.66$, $\beta_{1}=0.15$) and varied only a single parameter $\beta_{2}$ for comparison. The results in Table 7 reveal two critical findings: (1) CumPT achieves robust and stable performance with a single hyperparameter. Notably, the resulting Online Macro F1 score closely tracks the original Macro F1, indicating minimal performance degradation during the tiering process. (2) The traditional method exhibits high sensitivity to its hyperparameter settings, leading to significant variance in performance. This highlights a major practical challenge that identifying the optimal combination of its multiple hyperparameters requires extensive and costly experimentation.
@@ 4.4. Online Evaluation
To validate the effectiveness of the model in real online applications, we conducted a side-by-side evaluation based on manual annotation. We selected 2,000 queries, requested both experimental buckets for each query, and compared the top-10 results across three metrics. The specific results are shown in Table 8. We provide a fine-grained analysis of these results by segmenting the queries into four distinct categories. The most substantial gains were observed for queries seeking affordable alternatives. This is because such queries pose a classic challenge for traditional term-matching models, which often fail to distinguish between a target brand and its alternatives. For instance, a query for a “Miumiu alternative” would often cause a term-based model to incorrectly retrieve authentic Miumiu products with high relevance scores due to strong lexical overlap. Furthermore, significant improvements were also recorded across the other query categories. These results collectively demonstrate that the reasoning and general knowledge capabilities of our LLM-based model can effectively address long-standing challenges in search relevance—particularly for queries that require a deep, semantic understanding of user intent rather than simple lexical matching.

### v4

@@ 4.1.2. Baselines
We employ multiple models as our baselines, including BERT, Qwen3-0.6B, and Qwen3-30BA3B(Yang et al., 2025). Our model is iteratively developed based on Tbstar. Tbstar is a closed-source e-commerce large language model foundation trained from scratch by Taobao, utilizing over 1010T tokens of general corpus and rich domain-specific e-commerce corpus. Based on a Mixture of Experts (MoE) architecture, it has a total parameter count of 4242B with 3.53.5B activated parameters.
  • BERT: A 24-layer BERT model foundation trained from scratch based on general corpus and Taobao’s internal e-commerce corpus, subsequently fine-tuned on the full training set using formula 4.
@@
  • Qwen3-30BA3B: An open-source Mixture-of-Experts (MoE) model from the Qwen family, featuring 30 billion total parameters and 3 billion active parameters per forward pass. This model was also supervised fine-tunedon our full training set same as Qwen3-0.6B.
@@ 4.3.1. Ablation of Post-CoT
We contrast the “think-then-respond” paradigm with the “respond-then-think” paradigm (distinguished by the postCoT subscript) in Table 2. We evaluate the performance of each paradigm both at the SFT stage and subsequent DPO stage. At the SFT-only stage, the post-CoT approach demonstrates a significant advantage over the standard approach. This provides strong evidence that post-CoT effectively mitigates the issue of error accumulation in CoT, positioning it as a potentially superior approach for classification tasks. Furthermore, the application of DPO consistently yields notable performance gains under both configurations. This confirms that even within the post-CoT framework, there remains substantial headroom for improvement through reinforcement learning.
@@ 4.3.3. Ablation of CumPT
This section provides a detailed analysis of advantages of our CumPT method. Macro F1 is computed directly on model’s raw outputs, representing its intrinsic classification capability. Online Macro F1 is computed after the tiering mentioned in Sec. 3.4 (either the traditional formula-based approach via Eq.11, Eq.12 or CumPT). Additionally, we report Precision, Recall, and F1-score specifically for “Good” tier.
To illustrate the sensitivity of the traditional method, we fixed a subset of its many hyperparameters ($\alpha_{1}=0.33$, $\alpha_{2}=0.66$, $\beta_{1}=0.15$) and varied only a single parameter $\beta_{2}$ for comparison. The results in Table 7 reveal two critical findings: (1) CumPT achieves robust and stable performance with a single hyperparameter. Notably, the resulting Online Macro F1 score closely tracks the original Macro F1, indicating minimal performance degradation during the tiering process. (2) The traditional method exhibits high sensitivity to its hyperparameter settings, leading to significant variance in performance. This highlights a practical challenge that identifying the optimal combination of its multiple hyperparameters requires extensive and costly experimentation.
@@ 4.4. Online Evaluation
To validate the effectiveness of the model in real online applications, we first conducted a side-by-side evaluation based on manual annotation. We selected 2,000 queries, requested both experimental buckets for each query, and compared the top-10 results across three metrics. The specific results are shown in Table 8. We provide a fine-grained analysis of these results by segmenting the queries into four distinct categories. The most substantial gains were observed for queries seeking affordable alternatives. This is because such queries pose a classic challenge for traditional term-matching models, which often fail to distinguish between a target brand and its alternatives. For instance, a query for a “Miumiu alternative” would often cause a term-based model to incorrectly retrieve authentic Miumiu products with high relevance scores due to strong lexical overlap. Furthermore, significant improvements were also recorded across the other query categories. These results collectively demonstrate that the reasoning and general knowledge capabilities of our LLM-based model can effectively address long-standing challenges in search relevance –particularly for queries that require a deep, semantic understanding of user intent rather than simple lexical matching.
In addition, we conducted strict online A/B tests to evaluate changes in user purchase-related behaviors. The primary metrics include Item Page View (IPV), Transaction volume, and Gross Merchandise Value (GMV). The results show a 0.22% increase in UV, 2.43% increase in IPV, a 0.82% increase in Transaction volume, and a 0.29% decrease in GMV. Overall, these metrics remain essentially on par with those of the baseline bucket, indicating that the optimization of our relevance model does not impair users’ purchase intent while substantially improving their overall shopping experience.
@@ 5. Conclusion
Appendix A Appendix
A.1. Implementation Details
During the Supervised Fine-Tuning (SFT) phase, we used a batch size of $1024$ and a learning rate of 1e61e-6, managed by a cosine scheduler with a warmup ratio of $0.05$. We employed the AdamW (Loshchilov and Hutter, 2019) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The model was trained for $2$ epochs with a maximum input length of $4096$ tokens.
In the GRPO reinforcement learning phase, we configured a rollout batch size of $64$ with a learning rate of 1e61e-6. For each sample, we generated $16$ responses using a sampling temperature of $0.99$ and top-k of $100$. The thresholds for our online difficulty sampling were set to a range of $[0.01,0.9]$. For stabilization, we clipped the importance sampling ratio to a range of $1\pm 0.2$ and the advantage values to $\pm 2$. Our training pipeline is implemented based on the open-source framework ROLL(Wang et al., 2025).
A.2. Online Deployment
For online deployment, we directly served the TaoSR1 model in the production environment. Due to its relatively large size, the initial request Response Time (RT) was high, prompting us to implement several engineering optimizations to improve online throughput.
  • Optimizing load balance. Instead of using a fixed batch size, we partition incoming requests according to their sequence length, ensuring that the computational workload across batches is approximately balanced.
  • Scheduling across data-centers. We introduced proxy services for data centers in different regions, which poll and dispatch micro-batched requests to idle machines, thereby alleviating cross–data-center scheduling bottlenecks.
  • Refining model deployment. We enabled FP8 quantization for our relevance model, and refactored the key-value cache reuse mechanism to support prefix sharing within the same batch for requests with identical prefixes.
After these optimizations, the average online RT per user request decreased from 800 to 286 ms, and the Model FLOPs Utilization (MFU) finally reaches 46.2%. Taking advantage of the post-CoT generation paradigm adopted in our relevance model, the online system only needs to optimize latency in the prefill phase, without introducing additional overhead from the decoding phase.
A.3. Further Analysis of Post-CoT
To further analyze the effectiveness of Post-CoT and prevent it from overfitting to the offline evaluation set—which often leads to ”hacking” the metrics—we proposed a more objective evaluation method. Overfitting is a common issue; in our past experiments, we observed that simple tricks like swapping the query and item inputs or repeating the query could increase offline scores but failed to produce any real benefits in online deployment.
We aimed to build a benchmark that is independent of query and relevance grade distributions to objectively evaluate the Post-CoT model. Therefore, we established an Atomic Capability Evaluation Benchmarks (v1) for relevance modeling. This preliminary set allows us to derive initial insights. We divided the relevance judgment process into four dimensions: E-commerce General Knowledge, Business Rule Adherence, Reasoning, and Feature Utilization. Based on our analysis, most online relevance issues fall into one of these four categories.
We defined clear criteria for these capabilities and had annotators label a balanced dataset of 2,000 samples. We used Accuracy (Acc) to evaluate the following models:
[…]
Figure 2. Radar chart of model performance across Atomic Capability Evaluation Benchmarks.
Radar chart of model performance across Atomic Capability Evaluation Benchmarks
The results are shown in Figure 2. We draw several important conclusions:
  1. (1)
    Model Scale: More active parameters lead to overall improvements across all dimensions (shown by the Green Line). However, this creates greater challenges for online performance (e.g., latency).
  2. (2)
    E-commerce General Knowledge: Introducing RAG brings the most significant improvement in the dimension of E-commerce General Knowledge (shown by the Red Line).This demonstrates that RAG is a critical approach for supplementing the model’s missing knowledge.
  3. (3)
    Reasoning: Post-CoT does not essentially improve reasoning capabilities. Its gains come primarily from better adherence to business rules and more detailed utilization of features (comparing the Blue Line vs. the Grey Line).
Based on the current findings, it appears that significant improvements in reasoning capability may still rely heavily on the Pre-CoT paradigm. Consequently, implicit CoT or AdaptiveCoT may represent promising avenues for future exploration. Therefore, finding an optimal way to enhance online reasoning using CoT—while balancing model effectiveness, latency, and resources—remains a significant and challenging avenue for future research.